John is 2 years older than his sister Amber. Amber is only half of her father’s age.
If Amber’s father is 50 years old, how old is John?

Answer by TimothyLamb(4379)

j = a + 2
a = 50/2 = 25
---
j = 27
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php